Value of a stock is currently at $40. Volatility of that stock is 30% per year and risk-free interest rate with continuous compounding is at 5% per year. Suppose you are planning to value a 3-month European call option with strike price at $41 using a two-step binomial model. Answer the following using this information. (Binomial Tree Approach to Option Valuation describe how to solve this problem)

What are the values of u, d and q?
Answer #1

uret dre ovi = 1

u = e^(0.3*(3/12)^(1/2)) = 1.1618

d = 1/u = 1/1.1618 = 0.8607

Step 5 Binomial Tree Pricing Step 1 Step 2 Scenario price-Spot price*(U)^2=40*1.1618^2=53.99 Payoff HH=Max(Scenario price-Str